Date: Thu, 13 Nov 1997 11:38:05 -0800 (PST)
Subject: 2MASS WG Mtg #136 Minutes

           IPAC 2MASS Working Group Meeting #136 Minutes

Attendees: R. Beck, R. Cutri, T. Evans, J. Fowler, L. Fullmer,
           T. Jarrett, D. Kirkpatrick, G. Kopan, B. Nelson,
           B. Wheaton, S. Wheelock, J. White


1.) MPCAT Delivery
2.) MAPCOR Tuning
3.) PSF Model Status
4.) Science Team Meeting
5.) Telescope/Camera/Computer Status


1.) MPCAT Delivery

    J. Fowler reported that MPCAT 1.1 had been delivered. This
version includes computation of individual observation times
(i.e., correcting the scan midpoint time for each prediction's
declination distance from the midpoint, using the standard scan
rate of 58.9 arcsec/s), correction for proper motion that occurs
during the scan, clipping asteroid major-axis error bars at a
minimum of 2 arcsec, clipping 2MASS error bars at a maximum of
0.5 arcsec, and using a chi-square threshold of 16.0. The
acceptance region is therefore an ellipse with major and minor
axes of 8 and 4 arcsec, respectively, oriented as specified by
the predictor. The area is about 100 square arcsec. It may be
desirable to add code to the match-decision logic that uses the
MAPCOR status information.

2.) MAPCOR Tuning

    T. Evans, D. Kirkpatrick, and S. Wheelock reported that MAPCOR
parameter tuning had made great progress. The new parameter 
specifications had been delivered on 3 November; processing after
that date has been using these new values. Some aspects such as
bright-star processing still require further tuning activity, but
the overall status of artifact flagging is one of vast improvement.

3.) PSF Model Status

    B. Wheaton reported that the PSF model-generating code is
experiencing serious problems. Many IEEE error messages are being
produced. This is especially true when running K-band data, which
also experience much higher rejection rates than J or H. He is
implanting IEEE exception handlers and tracing the locations
where these errors occur. When the problems have been solved, it
is desirable to replace the "FIND" algorithm with a routine that
reads in the FREXAS detections; this should produce a speed gain
sufficient to allow the code to run on several calibration scans
within the pipeline for diagnostic purposes.

4.) Science Team Meeting

    R. Cutri reported on the highlights of last week's Science
Team meeting in Boston. The following is from his Project Meeting
report on this subject:

     Cutri, Beichman, Chester, and Mazzarella attended the 2MASS
     Science Team Meeting at CfA last week. High points included:
     1) Receipt of the ACT catalog which uses Astrographic
     Catalog positions mated to Tycho positions to improve proper
     motions for nearly 1 million Tycho astrometric reference
     stars; 2) Agreement to delete at least one set of galaxy
     magnitudes; 3) Some useful input on parameter tuning by a
     few Science Team Members. Low points included several
     attempts to recommend design changes to pipeline, which were
     explicitly forbidden unless compromises to the Level 1
     Requirements could be demonstrated.

5.) Telescope/Camera/Computer Status

    R. Cutri reported on the status of various hardware; the
following is from his Project Meeting report on this subject:
     Mt. Hopkins operations continue. Focus mechanism failed last
     week and one night was lost. Problem was found to only be a
    faulty connection that was repaired.

     Shipping damage to Southern telescope will be repaired at a
     machine shop in La Serena. Expected delay to start of ops to
     be only 3-4 weeks.

     Analysis of southern camera Ks array reveals a large number
     of noisy pixels that would result in numerous 'holes' in the
     coverage on the sky. Recommendation at the Science Team
     meeting was to replace the array now that the camera is
     open. An email from UMASS the day after the meeting stated
     that the array wouldn't be replaced until further lab tests
     would be made. As usual, there will be an attempt to try to
     minimize the problem in software.

     The northern tape I/O computer, an Ultra 2300, was delivered
     2 weeks ago. We are waiting for disks to arrive so that it
     can be put into operation.